OpenStudy (anonymous):

Mrs. Morgan is serving large cookies at a holiday party. She needs enough cookies for 19 children to get 2 cookies each. Each package of cookies contains 6 large cookies for $2.49. Mrs. Morgan bought 6 packages of cookies. Did she buy the right number of packages? Yes, 6 packages contain a great deal of cookies. Yes, the store only had 6 packages of cookies. No, Mrs. Morgan should know that some students may be able to eat 5 cookies. No, Mrs. Morgan should have bought 7 packages of cookies since she is two co

OpenStudy (anonymous):

Answer: No, Mrs. Morgan should have bought 7 packages of cookies since she is two co.. 19*2=38 38/6= 6.3333 Round up, 7 .

OpenStudy (anonymous):

ok

OpenStudy (phi):

There are 19 kids. If each gets one cookie, you need 19 cookies. If each gets 2 cookies, then you need 2*19= 38 cookies. How many cookies are in 6 packages? 6 packages x 6 cookies per pack= 36. Ooops! 36 is not enough.